A diver has a mass of 75 kg takes part in a competition by having to perform 10 dives off the 10m platform above the water. Before he takes part he snacks on an energy bar (150 g) which provides 15 kj of energy. How many would the diver need to consume to meet the necessary energy to perform all 10 of his dives?
1
Expert's answer
2012-02-23T10:00:55-0500
the necessary work:
W=10*10m*75kg*g(10m/s^2)=75000j
N=75000j/15000j=5
Answer: 5 (750g)